General Electric is an mp3 players’ producer. Demand and supply functions of mp3 players are as
follows:
QD= 1,450 - 25P (Demand)
QS= -100 + 75P (Supply)
Where P is the price of mp3 players.
1. Calculate the market equilibrium price/output combination.
2. What will happen in the market if the price of mp3 players increased to $40?
3. How will the market adjust?
4. What will happen in the market if the price of mp3 players increased to $10?
5. How will the market adjust?
1
Expert's answer
2020-04-23T11:51:18-0400

1.To find equilibrium price and quantity we need to solve this equation:

"1450-25P=-100+75P"

"100P=1550"

"P'=15.5"

"Q'=1450-25\times15.5"

"Q'=1062.5"

So, P=15.5 and Q=1062.5.

2.When the price is $40 the quantity demanded is 450, but the quantity supplied is 2900. So suplly exceedes suply.

3.In such circumstances the amount actually sold on the market will be 450. In long-run the suply curve will shift to the left and new equilibrium(with higher price and higher quantity) will be established.

4.When the price is $10, the quantity demanded is 1200, but the quantity suplied is 250.

5.The quantity which will be sold on the market will be 250. In long-run the suply curve will shift to the right and new equilibrium(with lower price and higher quantity) will be established.